OpenStudy (anonymous):

What is the d value of the following sequence? 3, -6, -15, -24, -33

OpenStudy (anonymous):

answer choices -7 -9 -3 *5

OpenStudy (amistre64):

what is the value that you have to add to the first term to get the second term?

OpenStudy (anonymous):

thats what i dont understand cause my teachers never taught me this my online class is teaching me this

OpenStudy (amistre64):

3, plus what value; equals -6? you surely know how to add by now

OpenStudy (amistre64):

3 + d = -6 -6 + d = -15 -15 + d = -24 -24 + d = -33 any of these will produce a solution to d. d is called the common difference of the arithmetic sequence